Circular No. 6738

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

V518 PERSEI = GRO J0422+32 A. Iyudin and F. Haberl, Max-Planck-Institut für Extraterrestrische

Physik, report: “V518 Per (= GRO J0422+32) was again spotted in two mini-outbursts starting on Mar. 24 (3.1 mCrab, significance > 5.1σ) and July 23 (3.3 mCrab, significance > 4.6σ). These mini-outbursts of V518 Per, together with that reported on IAUC 6605, confirmed the outburst period of ∼ 120 days, suspected on the basis of earlier x-ray and optical measurements (op.cit.). The periodical nova outbursts confirm similarities between V518 Per and the low-mass binary system Aql X-1 in the timing sequence and in the decrease of the peak-flux amplitude for the consecutive outbursts. The flux values during mini-outbursts were derived under the assumption of a Crab-like power-law spectrum. The data were analyzed from the publicly available ASM quick-look results (op.cit.). Observations of V518 Per at other wavelengths are strongly encouraged at the times of the suspected periodical outbursts.”

COMET C/1997 D1 (MUELLER) Ephemeris extension to IAUC 6626 (orbital elements MPC 30063):

1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 30 8^{h}27.^{m}62 +24^{o}59.^{′}9 3.067 2.300 34.^{o}0 14.^{o}2 12.1
Sept. 9 8 27.02 +23 20.2 2.907 2.279  43.1 17.6 11.9
19 8 24.79 +21 35.0 2.723 2.263  52.8 20.7 11.7
29 8 20.30 +19 40.9 2.520 2.252  63.2 23.4 11.5
Oct. 9 8 12.69 +17 33.3 2.302 2.248  74.3 25.3 11.3
19 8 00.77 +15 05.5 2.079 2.249  86.5 26.2 11.1
29 7 43.05 +12 08.4 1.861 2.257 100.0 25.7 10.9
Nov. 8 7 17.73 + 8 31.5 1.665 2.270 115.1 23.3 10.7
18 6 43.45 + 4 09.1 1.512 2.288 131.3 18.9 10.5
28 6 00.69 − 0 45.9 1.427 2.312 146.2 13.7 10.4
Dec. 8 5 13.39 − 5 31.0 1.428 2.342 151.6 11.6 10.5
18 4 28.15 − 9 15.8 1.518 2.376 142.6 14.6 10.7
28 3 50.34 −11 39.4 1.681 2.415 128.0 18.7 11.0

Total visual magnitude estimate by R. J. Bouma, Groningen, The Netherlands (0.25-m reflector): Sept. 11.13 UT, 12.6.

1997 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green